英語 (English)

発信型英語 (Communicative English) (医 ((放 B)2 年))

フェネリー, マーク · グレアム 2 単位前期月 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 This course aims to help students to develop their communicative language skills through activities and tasks that require real language use. The policy is to learn by doing and students will be asked to engage in real communication using the skills acquired.

【授業の概要】 Students are expected to take a positive attitude towards participation and preparation and will be required to work in pairs and small groups. Homework and preparation will be expected.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】

1. To help students to become confident communicators.

2. Develop active listening skills. 【授業の計画】

1. Introduction: self introductions

2. Shadowing in Conversation

3. Describing People

4. Mini Presentation 1: A Person I Admire

5. Active Listening

6. Talking about daily life

7. Describing Places

8. Reacting in Conversation

9. Mini Presentation 2: A Place I want to Visit

10. Giving Directions

11. Agreeing and Disagreeing

12. Talking About News and Events

13. Mini Presentation 3: My Important NEWS

14. Giving an Opinion

15. Mini Debate

16. Final Assessment 【教科書】 English Firsthand 1 (4th Edition) Student Book with CDs 【参考書等】 [参考資料] 【成績評価の方法】 Students will be assessed on class participation, homework and performance in mini-presentations.Attendance and participation: 45%, Homework: 15%, Mini-presentations/debate: 20%, Final Assessment 20%

【再試験の有無】原則として無し

【受講者へのメッセージ】 [注意]

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220894 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ フェネリー . (オフィスアワー: fenn@shikoku-u.ac.jp)

English

Communicative English (医 ((放 B)2 年)) Mark Fennelly·Graham 2 units 前期月 1 ・ 2

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ This course aims to help students to develop their communicative language skills through activities and tasks that require real language use. The policy is to learn by doing and students will be asked to engage in real communication using the skills acquired.

Outline⟩ Students are expected to take a positive attitude towards participation and preparation and will be required to work in pairs and small groups. Homework and preparation will be expected.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩

1. To help students to become confident communicators.

2. Develop active listening skills.

Schedule⟩

1. Introduction: self introductions

2. Shadowing in Conversation

3. Describing People

4. Mini Presentation 1: A Person I Admire

5. Active Listening

6. Talking about daily life

7. Describing Places

8. Reacting in Conversation

9. Mini Presentation 2: A Place I want to Visit

10. Giving Directions

11. Agreeing and Disagreeing

12. Talking About News and Events

13. Mini Presentation 3: My Important NEWS

14. Giving an Opinion

15. Mini Debate

16. Final Assessment Textbook⟩ English Firsthand 1 (4th Edition) Student Book with CDs Reference⟩ [参考資料] Evaluation Criteria⟩ Students will be assessed on class participation, homework and performance in mini-presentations.Attendance and participation: 45%, Homework: 15%, Mini-presentations/debate: 20%, Final Assessment 20%

Re-evaluation⟩ 原則として無し

Message⟩ [注意]

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220894 Contact (Office-Hour, Room, E-mail)⟩

⇒ Fennelly . (Office Hour: fenn@shikoku-u.ac.jp)